# Low-Light Video Enhancement via Spatial-Temporal Consistent Illumination and Reflection Decomposition

### v1

@@ Abstract
Low-Light Video Enhancement (LLVE) seeks to restore dynamic and static scenes plagued by severe invisibility and noise. One critical aspect is formulating a consistency constraint specifically for temporal-spatial illumination and appearance enhanced versions, a dimension overlooked in existing methods. In this paper, we present an innovative video Retinex-based decomposition strategy that operates without the need for explicit supervision to delineate illumination and reflectance components. We leverage dynamic cross-frame correspondences for intrinsic appearance and enforce a scene-level continuity constraint on the illumination field to yield satisfactory consistent decomposition results. To further ensure consistent decomposition, we introduce a dual-structure enhancement network featuring a novel cross-frame interaction mechanism. This mechanism can seamlessly integrate with encoder-decoder single-frame networks, incurring minimal additional parameter costs. By supervising different frames simultaneously, this network encourages them to exhibit matching decomposition features, thus achieving the desired temporal propagation. Extensive experiments are conducted on widely recognized LLVE benchmarks, covering diverse scenarios. Our framework consistently outperforms existing methods, establishing a new state-of-the-art (SOTA) performance.
@@ 1 Introduction
Low-light enhancement aims to enhance underexposed images and videos captured in low-light conditions Xu et al. (2022); Wang et al. (2021a), improving their visual quality while reducing noise. This is a longstanding and important topic in computer vision, with wide-ranging applications, such as improving portrait photography on mobile devices Ignatov et al. (2017); Hasinoff et al. (2016) and supporting various downstream tasks, including nighttime face recognition Ma et al. (2022); Wang et al. (2022a) and vehicle detection Fu et al. (2023a); Wu et al. (2022b).
@@
Figure 1: Our method consistently achieves SOTA performance on different LLVE datasets with the same network architecture.
Numerous methods have been proposed to enhance underexposed images using supervised learning Wang et al. (2021b); Xu et al. (2022, 2023b); Cai et al. (2023); Wang et al. (2022b), yielding impressive results. However, these methods still struggle to provide satisfactory performance when applied to videos, a popular media format. The key challenge when enhancing videos is the need for consistent enhancement results across corresponding locations in different frames, which vary both spatially and temporally. Current methods rely on paired training data, where the ground truth ensures this consistency. Unfortunately, there is a scarcity of real-world, high-quality, spatially aligned video pairs for dynamic scenes, resulting in limited generalization capabilities for unseen videos with varying conditions.
One solution is to incorporate parameterized enhancement models with physical significance into the enhancement process, reducing the reliance on training data, e.g., Retinex theory Land (1977). According to this theory, an observed image $I$ can be decomposed into its illumination component $L$ and reflectance component $R$ using the equation $I=L\otimes R$, where $\otimes$ denotes element-wise multiplication. Several approaches Zhang et al. (2019); Wu et al. (2022a); Fu et al. (2023b) have been developed to predict both $R$ and $L$ under normal illumination conditions for a low-light input image. However, a suitable strategy for obtaining the Retinex decomposition in normal-light conditions for low-light input videos is currently lacking. Ensuring both temporal and spatial consistency in decomposition is crucial, yet previous strategies were primarily designed for individual images, thus posing a challenge for video processing.
@@
In this paper, we introduce an innovative approach to attain a consistent Retinex decomposition for normal-light video outputs, ensuring consistency in both spatial and temporal dimensions. We leverage cross-frame correspondences and real-world physical continuity constraints to achieve this decomposition without the need for explicit supervision. Furthermore, to enhance the consistency of the decomposition, we’ve developed an efficient dual-structure enhancement network featuring a novel cross-frame interaction mechanism.
Differing from some previous methods Wang et al. (2019a, 2021a), which directly assume the reflection term as the enhanced version, our approach follows Cai et al. (2017); Wei et al. (2018); Wu et al. (2022a); Fu et al. (2023b), i.e., aims to recover both the reflection and illumination components of normal-light images from the provided low-light inputs. Previous methods have verified that satisfying image-level reflection and illumination decomposition can be achieved with suitable reflection and illumination priors via a data-driven manner. In this paper, we aim to design appropriate priors for video data to implement the decomposition.
@@
We recognize that the reflectance component $R$ should represent the physical properties of objects, which should remain consistent across various observation perspectives Wei et al. (2018). To achieve this, we first calculate corresponding locations in videos and utilize these computed spatial-temporal correspondences, along with uncertainty values, to link the reflection terms across different frames. Conversely, the illumination terms are subject to a spatial smoothness loss to model real-world physical constraints Wei et al. (2018). This approach allows us to establish a consistency constraint specifically for temporal-spatial illumination and appearance in the enhanced videos, without requiring additional supervision information.
Additionally, the architecture of the enhancement networks plays a crucial role in shaping the decomposition results. It is important that different frames share features and receive simultaneous supervision, promoting the alignment of feature representations to synthesize Retinex parameters. In our paper, we introduce a novel dual network structure capable of achieving such feature propagation. Two frames are used as inputs, and their features are shared within the designed Cross-Frame Interaction Module (CFIM). CFIM facilitates interaction through a combination of long-range cross-frame attention computation and short-range channel-spatial fusion, ensuring the accurate and comprehensive merging of features from both global and local perspectives. CFIM can be seamlessly integrated into any encoder-decoder single-image enhancement structure by inserting it at the deepest hidden layer. In this paper, we illustrate this with the simplest U-Net Ronneberger et al. (2015) as an example, and it’s worth noting that CFIM comprises only 0.4M parameters.
@@
We performed experiments on various widely recognized video enhancement datasets Wang et al. (2021a); Chen et al. (2019); Fu et al. (2023a); Zhang et al. (2021). Our results, both quantitative and qualitative, consistently demonstrate the effectiveness and state-of-the-art (SOTA) performance of our framework, as shown in Fig. 1. Furthermore, we conducted a large-scale user study involving 100 participants, which clearly showcased the superiority of our enhanced results in terms of human subjective perceptions.
In summary, our contributions are three-fold.
@@
  • We introduce a novel canonical form for LLVE that predicts spatial-temporal consistent Retinex decomposition for normal-light outputs. This innovative decomposition strategy leverages spatial-temporal correspondences and continuity.
@@
  • We design a new LLVE network that facilitates interaction among the features of different frames, leading to a consistent decomposition.
@@ 2 Related Work
Low-Light Image Enhancement. To enhance the quality of a low-light video, image enhancement methods can be applied on a frame-by-frame basis. In recent years, learning-based Low-Light Image Enhancement (LLIE) techniques Yan et al. (2014, 2016); Lore et al. (2017); Cai et al. (2018); Zamir et al. (2020); Xu et al. (2020); Zeng et al. (2020); Kim et al. (2021); Zhao et al. (2021); Zheng et al. (2021); Wang et al. (2021b); Liu et al. (2021); Yang et al. (2021b); Jiang et al. (2021); Yang et al. (2021a) have made significant advancements, with a primary focus on supervised approaches thanks to the abundance of image pairs available for training. Xu et.al. Xu et al. (2022) introduced spatial-varying operations, considering Signal-to-Noise Ratio (SNR) as a prior factor. However, when applying image enhancement algorithms to individual frames, the issue of flickering often arises.
Low-Light Video Enhancement. In addition to the need for LLIE, there is a growing demand for video enhancement, considering the widespread use of videos as a popular data format on the internet and in photographic equipment. Various approaches have been proposed in this context Chen et al. (2019); Triantafyllidou et al. (2020); Ye et al. (2023); Lv et al. (2023); Xu et al. (2023a); Fu et al. (2023a). Danai et.al. Triantafyllidou et al. (2020) introduced a data synthesis mechanism to generate dynamic video pairs from SID Chen et al. (2019). Wang et.al. Wang et al. (2021a) proposed a multi-branch network that simultaneously estimates noise and illumination, suitable for videos with severe noise conditions. Liu et.al. Liu et al. (2023) and Liang et.al. Liang et al. (2023) used prior event information to learn enhancement mapping for brightening videos. Xu et.al. Xu et al. (2023a) designed a parametric 3D filter tailored for enhancing and sharpening low-light videos. Recently, Fu et.al. Fu et al. (2023a) introduced a video enhancement method called LAN, which iteratively refines illumination and adaptively adjusts it. However, it’s important to note that LAN lacks an explicit constraint for maintaining consistent reflectance and illumination decomposition.
Several datasets for video enhancement, encompassing both static Chen et al. (2019); Jiang and Zheng (2019); Wang et al. (2019b); Triantafyllidou et al. (2020) and dynamic motions Wang et al. (2021a); Fu et al. (2023a), have been introduced. In this paper, we introduce a novel LLVE method, designed to enhance effects on these datasets. Our decomposition method explicitly and consistently models reflectance and illumination for all frames, with the guidance of the videos themselves.
@@ 3 Method
We commence by revisiting the formulation of the Retinex theory in Sec. 3.1. Following that, we provide a detailed exposition of our decomposition strategy specifically tailored for video data in Sec. 3.2. Subsequently, in Sec. 3.3, we delve into the description of our cross-frame interaction. Finally, the training loss for our framework will be introduced in detail.
Figure 2: Our framework provides a comprehensive solution that includes explicit and consistent modeling for Retinex decomposition of enhanced normal-light outputs across different frames. To achieve this, we enforce consistent features in the reflectance terms across different frames by leveraging computed correspondences in the temporal dimension of videos ($\mathcal{L}(R_{t})$). Simultaneously, we ensure that the illumination terms exhibit a spatially continuous distribution ($\mathcal{L}(L_{t})$), aligning with real-world scenarios. Furthermore, our network incorporates cross-frame interaction and simultaneous supervision of different frames within a video, encouraging consistent Retinex features among these frames. For a more detailed visual representation, please refer to Fig. 3.
3.1 Retinex Model
Retinex Theory. When presented with an image $I$ (in this paper, we denote the low-light image as $I_{d}$ and the normal-light image as $I_{n}$, with $d$ and $n$ serving as subscript abbreviations), we assume its decomposition into illumination $L$ and reflectance $R$ follows the Retinex theory Land and McCann (1971); Fu et al. (2023b), as
$I=L\otimes R,$ (1)
@@ 3.1 Retinex Model
where $\otimes$ denotes the element-wise multiplication, $L$ and $R$ can be formulated for different channels Wang et al. (2019a); Cai et al. (2023), i.e., the channel number is 3 for the sRGB domain. In this context, $L$ describes the light intensity of objects, which is expected to exhibit piece-wise continuity. On the other hand, $R$ represents the physical properties of the objects, encompassing textures and details observed in the data $I$. If we obtain $L$ and $R$ in the normal-light conditions for a given input low-light image, the target normal-light image can be acquired accordingly. Numerous methods have been proposed to address this decomposition problem Zhang et al. (2019); Wu et al. (2022a); Fu et al. (2023b), and their common objective can be summarized as
$\mathcal{L}=\|I-L\otimes R\|+\mathcal{L}(L)+\mathcal{L}(R),$ (2)
where $\mathcal{L}(L)$ and $\mathcal{L}(R)$ denote the constraints for illumination $L$ and reflectance $R$, respectively.
Shortcoming of current Retinex models for videos. Several prior approaches have been devised to attain this desired decomposition by imposing different constraints on the two components, such as structural and smoothness constraints Li et al. (2018); Wei et al. (2018). Nevertheless, there is a scarcity of satisfactory Retinex decomposition prediction tailored for sequential video data. The issue concerning the decomposition of video data can be summarized as
$\mathcal{L}_{v}=\mathbb{E}_{t=1:T}[\mathcal{L}_{R,t}+\mathcal{L}(L_{t})+\mathcal{L}(R_{t})],\quad\mathcal{L}_{R,t}=\|I_{t}-L_{t}\otimes R_{t}\|,$ (3)
@@
where $\mathbb{E}$ is the average operation, $T$ is the number of frame. Compared with Eq. 2, we need to simultaneously guarantee the decomposition quality and the consistency among different frames’ results.
@@
Video data guide Retinex models by themselves. We have discovered that video data itself can play a crucial role in facilitating the achievement of a satisfactory Retinex decomposition prediction for normal-light outputs, thus obviating the need for external priors typically used in low-light enhancement. By establishing correspondences among different frames, we can impose specific constraints: the reflectance of each frame, denoted as $R_{t}$, should faithfully represent the intrinsic texture of objects within the target scene, remaining consistent regardless of changes in viewpoint. Likewise, the illumination of each frame, denoted as $L_{t}$, should exhibit the desired continuity consistent with real-world physical properties. For further elaboration, please refer to Sec. 3.2.
@@
Problem formulation. We adopt the supervised setting. Given a clip of low-light data $I_{d,t},t\in[1,T]$, there is a paired normal-light data $I_{n,t},t\in[1,T]$. We aim to directly obtain the decomposition of $I_{n,t}$ from $I_{d,t}$ using network $f$, as
$\hat{L}_{n,t},\hat{R}_{n,t}=f(I_{d,t}),\hat{I}_{n,t}=\hat{L}_{n,t}\otimes\hat{R}_{n,t},$ (4)
where $\hat{L}_{n,t}$ and $\hat{R}_{n,t}$ are the estimated targets (several examples are displayed in Fig. 4), and $\hat{I}_{n,t}$ is the predicted enhancement result.
Figure 3: The lightweight cross-frame interaction mechanism (CFIM) to propagate different frames’ features, along with the cross-frame attention and spatial-channel fusion. Cross-frame interaction can be employed in the deep feature space of arbitrary single-image encoder-decoder frameworks, and we choose U-Net here.

### v2

# Low-Light Video Enhancement via Spatial-Temporal Consistent Decomposition

@@ Abstract
Low-Light Video Enhancement (LLVE) seeks to restore dynamic or static scenes plagued by severe invisibility and noise. In this paper, we present an innovative video decomposition strategy that incorporates view-independent and view-dependent components to enhance the performance of LLVE. We leverage dynamic cross-frame correspondences for the view-independent term (which primarily captures intrinsic appearance) and impose a scene-level continuity constraint on the view-dependent term (which mainly describes the shading condition) to achieve consistent and satisfactory decomposition results. To further ensure consistent decomposition, we introduce a dual-structure enhancement network featuring a cross-frame interaction mechanism. By supervising different frames simultaneously, this network encourages them to exhibit matching decomposition features. This mechanism can seamlessly integrate with encoder-decoder single-frame networks, incurring minimal additional parameter costs. Extensive experiments are conducted on widely recognized LLVE benchmarks, covering diverse scenarios. Our framework consistently outperforms existing methods, establishing a new SOTA performance.
@@ 1 Introduction
Low-light enhancement aims to enhance underexposed images and videos captured in low-light conditions (Xu et al., 2022; Wang et al., 2021), improving their visual quality while reducing noise. This technique can be applied in wide-ranging applications, such as portrait photography on mobile devices (Ignatov et al., 2017; Hasinoff et al., 2016), nighttime face recognition (Ma et al., 2022; Wang et al., 2022) and vehicle detection (Fu et al., 2023a; Wu et al., 2022). The key challenge when enhancing videos is the need for consistent enhancement results across corresponding locations in different frames, which vary both spatially and temporally. Moreover, there is a scarcity of real-world, high-quality, spatially aligned video pairs for dynamic scenes, resulting in limited generalization capabilities for unseen videos with varying conditions.
@@
Figure 1: Our proposed LLVE method consistently achieves SOTA performance on different LLVE datasets involving various scenes with the same network architecture.
One solution is to incorporate parameterized enhancement models with physical significance into the enhancement process, reducing the reliance on training data. According to the linear Lambertian model for intrinsic image decomposition (Yuille, 2012; Narihira et al., 2015), an observed image $\boldsymbol{I}$ can be expressed as the element-wise product ($\otimes$) of its albedo component and shading component. The albedo part is normally view-independent, capturing the intrinsic appearance, whereas the shading component is commonly view-dependent, influenced by lighting conditions and the surface normal’s direction. Inspired by the Lambertian model, we propose that a frame in the LLVE task can similarly be decomposed as $\boldsymbol{I}=\boldsymbol{L}\otimes\boldsymbol{R}$, where $\boldsymbol{L}$ represents the view-dependent term, and $\boldsymbol{R}$ corresponds to the view-independent component.
@@
In this paper, we introduce an innovative approach to attain our decomposition for normal-light video outputs, ensuring consistency in both spatial and temporal dimensions. We leverage cross-frame correspondences and real-world physical continuity constraints to achieve this decomposition without the need for explicit supervision. Furthermore, to enhance the consistency of the decomposition, we’ve developed an efficient dual-structure enhancement network featuring a novel cross-frame interaction mechanism.
@@
Our approach aims to predict both the $\boldsymbol{R}$ and $\boldsymbol{L}$ of normal-light images from the provided low-light inputs. It is important to note that the decomposition of view-independent and view-dependent terms is not unique, and our goal is to implement a suitable decomposition that enhances the performance of LLVE. In this paper, we aim to design appropriate priors directly from the video data to implement the decomposition. We recognize that $\boldsymbol{R}$ should represent the physical properties of objects, which should remain consistent across various observation perspectives. To achieve this, we first calculate corresponding locations in videos and utilize these computed spatial-temporal correspondences, along with uncertainty values, to link the predictions of $\boldsymbol{R}$ across different frames. Conversely, the terms of $\boldsymbol{L}$ are subject to a spatial smoothness loss to model real-world physical constraints (since the lighting field is continuous in practice). This approach allows us to establish a temporal-spatial consistency constraint in the enhanced videos, without requiring additional supervision. Note that we cannot guarantee the decomposition results are strictly view-independent or view-dependent, as this is not a 3D method. Our goal is to apply appropriate constraints to approximate desired decomposition properties, while focusing on optimizing the final LLVE performance with our approach.
Additionally, we introduce a dual-network structure to enforce the consistency of feature representations across frames, facilitating the consistent synthesis of decomposition in turn. Two frames are used as inputs, and their features are propagated within the designed Cross-Frame Interaction Module (CFIM). CFIM differs from other similar architectures in the restoration task (Lv et al., 2023) in terms of the fusion manner. CFIM facilitates interaction through a combination of long-range cross-frame attention computation and short-range channel-spatial fusion, ensuring the accurate and comprehensive merging of features from both global and local perspectives. Moreover, CFIM selectively incorporates knowledge from randomly chosen neighboring frames into the backbone (i.e., the part without feature propagation) during training. In this way, CFIM help the backbone in learning how to adaptively utilize knowledge from various cross-frame features, resulting in a more robust backbone.
@@
We performed experiments on various widely recognized video enhancement datasets. Our results, both quantitative and qualitative, consistently demonstrate the effectiveness and state-of-the-art (SOTA) performance of our framework, as shown in Fig. 1. Furthermore, we conducted a large-scale user study involving 100 participants, which showcased the superiority of our results in terms of human subjective perceptions. In summary, our contributions are three-fold.
@@
  • We introduce a novel canonical form for LLVE that predicts spatial-temporal consistent decomposition with view-independent and view-dependent terms for normal-light outputs. This decomposition strategy leverages spatial-temporal correspondences and continuity.
@@
  • We design a new LLVE network that facilitates interaction among the features of different frames and fits our consistent decomposition strategy.
@@ 2 Related Work
Low-Light Image Enhancement. To enhance the quality of a low-light video, image enhancement methods can be applied on a frame-by-frame basis. In recent years, learning-based Low-Light Image Enhancement (LLIE) techniques (Jiang et al., 2021; Yang et al., 2021) have made significant advancements, with a primary focus on supervised approaches.
Low-Light Video Enhancement. In addition to the need for LLIE, there is a growing demand for video enhancement, considering the widespread use of videos as a popular data format on the internet and in photographic equipment. Various approaches have been proposed in this context (Chen et al., 2019; Triantafyllidou et al., 2020; Ye et al., 2023; Lv et al., 2023; Xu et al., 2023a; Fu et al., 2023a). Liu et.al. (Liu et al., 2023) and Liang et.al. (Liang et al., 2023) used prior event information to learn enhancement mapping for brightening videos. Xu et.al. (Xu et al., 2023a) designed a parametric 3D filter tailored for enhancing and sharpening low-light videos. Recently, Fu et.al. (Fu et al., 2023a) introduced a video enhancement method called LAN, which iteratively refines illumination and adaptively adjusts it. However, it’s important to note that LAN lacks an explicit constraint for maintaining consistent reflectance and illumination decomposition.
Several datasets for video enhancement, encompassing both static (Chen et al., 2019; Jiang and Zheng, 2019; Wang et al., 2019; Triantafyllidou et al., 2020) and dynamic motions (Wang et al., 2021; Fu et al., 2023a), have been introduced. In this paper, we introduce a novel LLVE method, designed to enhance effects on these datasets. Our decomposition method explicitly and consistently models view-dependent and -independent for all frames.
Figure 2: Our framework offers a comprehensive solution that explicitly and consistently models the view-independent and view-dependent decomposition of enhanced normal-light outputs across different frames. To achieve this, we enforce consistent features in the view-independent terms across different frames by leveraging computed correspondences in the temporal dimension of videos ($\mathcal{O}(\boldsymbol{R}_{t})$). Simultaneously, we ensure that the view-dependent terms exhibit a spatially continuous distribution ($\mathcal{O}(\boldsymbol{L}_{t})$), aligning with real-world scenarios. Furthermore, our network incorporates cross-frame interaction and simultaneous supervision of different frames within a video, encouraging consistent features for these frames derived from one video. For a more detailed visual representation, please refer to Fig. 3.
@@ 3 Method
3.1 Decomposition Model
Motivation. According to the linear Lambertian model (Narihira et al., 2015), an observed image $\boldsymbol{I}$ can be formulated as the element-wise product of its albedo and shading component. Here, the albedo part is view-independent, capturing the intrinsic appearance, while the shading component is view-dependent. Inspired by this, when presented with an image $\boldsymbol{I}$ (in this paper, we denote the low-light image as $\boldsymbol{I}_{d}$ and the normal-light image as $\boldsymbol{I}_{n}$, with $d$ and $n$ serving as subscript abbreviations), we assume its decomposition into view-dependent part $\boldsymbol{L}$ and view-independent part $\boldsymbol{R}$, as
$\boldsymbol{I}=\boldsymbol{L}\otimes\boldsymbol{R},$ (1)
@@ 3.1 Decomposition Model
where $\otimes$ denotes the element-wise multiplication, $\boldsymbol{L}$ and $\boldsymbol{R}$ can be formulated for different channels, i.e., the channel number is 3 for the sRGB domain. In this context, $\boldsymbol{L}$ mainly describes the light intensity of objects, which is expected to exhibit piece-wise continuity. On the other hand, $\boldsymbol{R}$ primarily represents the physical properties of the objects, encompassing textures and details observed in the data $\boldsymbol{I}$. If we obtain $\boldsymbol{L}$ and $\boldsymbol{R}$ in the normal-light conditions for a given input low-light image, the target normal-light image can be acquired accordingly. The common objective can be summarized as
$\mathcal{O}=\|\boldsymbol{I}-\boldsymbol{L}\otimes\boldsymbol{R}\|+\mathcal{O}(\boldsymbol{L})+\mathcal{O}(\boldsymbol{R}),$ (2)
where $\mathcal{O}(\boldsymbol{L})$ and $\mathcal{O}(\boldsymbol{R})$ denote the constraints for $\boldsymbol{L}$ and $\boldsymbol{R}$ (we will introduce our designed new constraints in Sec. 3.2). When applying this model to sequential video data, the objective related to the decomposition of video data can be summarized as
$\mathcal{O}_{v}=\mathbb{E}_{t=1:T}[\mathcal{O}_{rec,t}+\mathcal{O}(\boldsymbol{L}_{t})+\mathcal{O}(\boldsymbol{R}_{t})],$ (3)
$\mathcal{O}_{rec,t}=\|\boldsymbol{I}_{t}-\boldsymbol{L}_{t}\otimes\boldsymbol{R}_{t}\|,$
@@
where $\mathbb{E}$ is the average operation, $T$ is the number of frame, and $\boldsymbol{L}_{t}$/$\boldsymbol{R}_{t}$ is the decomposition output at the time index of $t$. Compared with Equation 2, we need to simultaneously guarantee the decomposition quality and the consistency among different frames’ results.
@@
Video Data Guide the Decomposition by Themselves. We have discovered that video data itself can play a crucial role in facilitating the achievement of our decomposition predictions for normal-light outputs, thus obviating the need for external priors typically used in low-light enhancement. By establishing correspondences among different frames, we can impose specific constraints: the $\boldsymbol{R}_{t}$ of each frame should faithfully represent the intrinsic texture of objects within the target scene, remaining consistent regardless of changes in viewpoint. Likewise, the $\boldsymbol{L}_{t}$ of each frame should exhibit the desired continuity consistent with real-world physical properties. For further elaboration, please refer to Section 3.2.
@@
Problem Formulation. We adopt the supervised setting. Given a clip of low-light data $\boldsymbol{I}_{d,t},t\in[1,T]$, there is a paired normal-light data $\boldsymbol{I}_{n,t},t\in[1,T]$. We aim to directly obtain the decomposition of $\boldsymbol{I}_{n,t}$ from $\boldsymbol{I}_{d,t}$ using network $f$, as
$\hat{\boldsymbol{L}}_{n,t},\hat{\boldsymbol{R}}_{n,t}=f(\boldsymbol{I}_{d,t}),\hat{\boldsymbol{I}}_{n,t}=\hat{\boldsymbol{L}}_{n,t}\otimes\hat{\boldsymbol{R}}_{n,t},$ (4)
where $\hat{\boldsymbol{L}}_{n,t}$ and $\hat{\boldsymbol{R}}_{n,t}$ are the estimated targets, and $\hat{\boldsymbol{I}}_{n,t}$ is the predicted enhancement result. Our framework is shown in Fig. 2 that will be introduced in the following sections.